Grade 11Physical Chemistry

Calculate the energy emitted when electrons of 1g atom of hydrogen undergo transition giving the spectral lines of lowest energy in the visible region of its atomic spectra.

Srinu Matala
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Ans: E =hc/λ
1/λ  = R[1/n1- 1/n22]
=R[1/22  -1/32]
=(109678)*(5/36)
1/λ  = 15233.055555555555555555555555556
λ =6.7*10- cm
E =hc/λ
= (6.625*10-27)*(3*1010)/ (6.7*10-5)
=(2.97)*10-12 ergs
Energy 1 Hydrogen = 2.97 *10-12 ergs
Energy emitted from 1gramof  Hydrogen =(2.97 *10-12 ergs)*Avogadronumber
=(2.97 *10-12 ergs)*6.02*1023
= 17.8794*109 ergs
= 17.88*106 k Ergs